Please HELP.
PLEASE DO ALL THANK YOU (:
4. In a certain human population, the genotypes of the M and N blood groups are to in the following frequencies: 179 a. What are the frequencies of the M and N alleles? b. According to the Hardy-Weinberg principle. what are the expected frequencies of the three genotypes? Are the expected and observed genotype frequencies similar or different? Suggest a hypothess to explain this observation. c. 5. Imagine that you have a population with a gene that has two alleles, X and Y. The frequency of X is 60% and Y is 40% Genotype frequencies are XX = 20% , 60% and YY-20% Is the population is Hardy Weinberg Equilibrium? Please show all your work.
Answer #1

4. a) Frequency of allele M (p)= [(2*genotype frequency of MM)+(1*genotype frequency of MM)]/(2*Total frequency)

= [(2*0.179)+(1*0.502)]/(2*1)=0.43

Frequency of allele m (q)= [(2*genotype frequency of mm)+(1*genotype frequency of MM)]/(2*Total frequency)

= [(2*0.319)+(1*0.502)]/(2*1)=0.57

b) According to Hardy-Weinberg equation,

p2+2pq+q2=1

0.432+(2*0.43*0.57)+0.572=1

So, this population is in Hardy-Weinberg equillibrium.

c) Expected frequency of MM =0.432=0.1849

Expected frequency of Mm =2*0.43*0.57=0.4902

Expected frequency of mm =0.572=0.3249

The expected and observed frequencies are different. This difference may be due to mutations or selection which changes the genotype frequency

5. p = [(2*20)+(1*60)]/(2*100)=0.5

q=[(2*20)+(1*60)]/(2*100)=0.5

p2=0.52 = 0.25

q2=0.52 = 0.25

2pq=2*0.5*0.5=0.5

These expected genotype frequencies are not equal to that observed. So they are not in equillibrium.